OpenStudy (anonymous):

help please? f (x )= -x^2+4x+2 the minimum y-value is y= @HawkCrimson

OpenStudy (anonymous):

okay, so you want to minimize this function. Are you in calculus?

OpenStudy (anonymous):

algebra 2

OpenStudy (anonymous):

Okay, so you need to find the vertex of the parabola. Do you know how to do that?

OpenStudy (anonymous):

wait hold on, -x^2?

OpenStudy (anonymous):

Are you sure it's not maximum?

OpenStudy (anonymous):

yes -x^2

OpenStudy (anonymous):

sorry, yes, maximum, my mistake

OpenStudy (anonymous):

-x^2+4x+2=6-(x-2)^2, now since (x-2)^2 is always nonnegative, ....

OpenStudy (anonymous):

I come up with -2... but is this right?

OpenStudy (anonymous):

Ah, that makes more sense. So, we solve for the x value that gives us the highest value, and then plug it into our equation.

OpenStudy (anonymous):

Should be 2

OpenStudy (anonymous):

Then plug it into your equation to get 6

OpenStudy (anonymous):

can you show me from 2?

OpenStudy (anonymous):

so, the +2 at the end is just a vertical shift, so it doesn't matter for finding the maximum x value. So, we just have the equation -x^2 + 4x = 0, and find the zeros. This gives us the two zeros, and the vertex by definition, is halfway between them. The equation is x(4-x)=0 so our zeros are x=0 and x=4. taking the average value (4+0)/2 = 2 Plugging this number back into our equation, we get -(2)^2 + 4(2) + 2 = -4+8+2 = 6

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

ah, I see where I messed up the negative. so 6 is my maximum.

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

thanks for the explanation :)